OpenStudy (anonymous):

Functions f(x) and g(x) are shown below: f(x) = 2 tan(3x + π) g(x) = 3 sin(4x - π) - 2 Using complete sentences, explain how to find the y-intercept for each function and determine which function has the largest y-intercept

OpenStudy (anonymous):

someone please help meeee :c

OpenStudy (sidsiddhartha):

put x=0 f(0) = 2 tan(3(0) + π) =2 tan(pi) similarly for the other function and check which is greater

OpenStudy (anonymous):

Thank you so much but I'm literally clueless when it comes to trig, can you tell me what the other function would equal and which is greater? pleeease <333

OpenStudy (sidsiddhartha):

for the second function g(0)=3sin(-pi)-2=-3sin(pi)-2=-2.16 and f(0)=0.109 so f(0)>g(0)

OpenStudy (anonymous):

thank youuuuuuu I love you